bill533:

Which of the following sequences is an arithmetic sequence? A. 1, 4, 16, 64, ... B. 75, 70, 65, 60, ... C. 1600, 400, 100, 25, ... D. 6, 12, 24, 48, ...

bill533:

@dude

bill533:

@AngeI

dude:

Arithmetic sequences are usually adding or subtracting to get the next/previous number

bill533:

an arithmetic sequence always has the same difference between successive terms That is (term two - term one) = ( term three - term two) = ( term four - term three) ......etc.... Which one of these fit the bill ???

dude:

Yeah

bill533:

i'm thinking D ?

bill533:

actually wait

bill533:

i think it's A

dude:

You just stated the rule used Subtract the first and second 4-1 = 3 Second and third 16-4 = 12 They're not equal

bill533:

YEs you're right

bill533:

i believe it's B

dude:

Good

bill533:

COULD YOU SHOW A DEMONSTRATION FOR B

bill533:

how they're all equal

dude:

I mean you should have gotten it on your own but First and second 70-75= -5 Second and Third 65-70 = -5 Third and fourth 60-65 = -5

bill533:

ALRIGHT BECAUSE I WANTED TO SEE IT YOUR WAY BUT I SEE

bill533:

thanks